SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Trigger Configuration History

Definitions of L1 trigger objects
Instructions for retrieving TrgConfig from DB

Level-1 Physics Configuration 

Start 
Run
Start 
Time
Configuration 
name
Key  Comments
64234
Apr-26-2006
PHYSICS
0x1b7d The "PHYSICS_TIGHT" configuration tested last week is now used in the default "PHYSICS" key.
63994
Apr-18-2006
PHYSICS_TIGHT
0x1b21
Similar to the "PHYSICS" configuration, except that 1M was added to 2Zt&1A and the prescale on 1Z was increased from 64 to 512
62694-
Mar-3-2006
PHYSICS_BURST
0x1a3c
The extra events will be logged on their own "L3OutRadiationSample" line. This runtype temporarily replaces our standard HER/LER injection samples with the new SVT radiation sample.
56797
Aug-3-2005

0x1653/1652
First run with the DCH half waveform bug fixed. The PHYSICS_DCHWF key was used, but then changed back to PHYSICS after confirming the bug was fixed.
53726
Apr-16-2005

0x1524
First colliding beam run of Run 5.
50636
Jul-31-2004
Physics_Ztrigger
0x1359
First DCZ physics configuration.
46477
Mar-25-2004

0xce7
First colliding beam run after IR vacuum fix.
45992
Mar-12-2004

0xcab
Start of regular HER/LER double trickle.
42798
Dec-04-2003

0xbf4
Start of regular LER trickle.
40003
Sep-09-2003

0x989
First colliding beam run of Run 4.
- Apr-20 2003 - - The emc was unbiased from the time of the first power cut at 9:51pm Sunday 4/20 until 9:22am Monday 4/21
32191 Nov-14 2002 Physics-61 - EMC frontend energy scale constants updated.
- Nov-07 2002 - - From the begining of run 3 till Nov 7 DIOM is sector 1 caused reframing problem.
29113 June-18-2002 Physics-61 1445 Renamed PHYSICS_TEMP to PHYSICS. So no towers are masked off and the EMC noisy crystals are still masked off.
29017 June-14-2002 Physics-61 1437 Unmasked tower in (0,3). Run type is PHYSICS_TEMP. EMC noisy crystals still masked off.
28831 June-06-2002 Physics-61 1422 Masked off hot tower tower in (theta,phi)=(0,3). In runs 28931, 28933 and 28934 the EMC had 4+3 noisy crystals in the endcap (theta~2, phi~30 and theta=4, phi~40 in EMC JAS plot p3) affecting tower (0,15). M occupancy was down in that region. These crystals were masked off from run 28935. Note that the configuration key did not change.
28765 June-03-2002 Physics-61 1410 Unmasked tower in (theta,phi)=(0,3) as calibration showed it to be OK. Hot again at beginning of run 28828 (at least).
28671 May-30-2002 Physics-61 1374 Masked off hot tower in (theta,phi)=(0,3).
27650 April-15-2002 Physics-61 1315 Unmasked hot tower in (theta,phi)=(0,3) as calibration showed it was OK. Alias is PHYSICS. It came back in run 28669.
27631 April-14-2002 Physics-61 1330 Masked off hot tower in (theta,phi)=(0,3). Alias is PHYSICS_EMT_HOTTOWER.
27390 April-5-2002 Physics-61 1312 Unmasked EMT tower (calibration was OK). Alias is PHYSICS_EMT. New hot tower appeared in (theta,phi)=(0,3) in run 27622.
27326 April-2-2002 Physics-61 1298 EMT map 133. Masked off hot tower in (theta,phi)=(0,17) (counting from 0). Alias is PHYSICS_DCH_DEBUG (been used for some time now).
27028
Mar-22-2002
Physics-61 1278
IFT barrel majority OR logic changed from 4/8 to 3/8 to increase efficiency. Endcap still 4/8.
25871 February-4-2002 Physics-61 1265 Same as previous - Physics_emcwf, but right trgconfig. Everything unmasked. The 'bunch crossing' signal broke around January 28th. It stayed broken until we pulled the plug on February 10th (before the background studies). CPO put the ringturn signal back in on February 14 - 136kHz. Spike in SL5 from run 27118 to 27130. Warm tower in (theta,phi)=(0,17) (counting from 0) from 27308 to 27310. Run 27311 is OK, and then it's hot from 27312
25734 January-30-2002 Physics-60 ! 1251 Physics_emcwf_nohot with EMT map 136: Swapped UPCs. Everything unmasked. Only (0,13) is warm. Later (January 31) the same map is used for physics_emcwf. From run 25777 (0,14) becomes warm again. Then Stefania touches the (0,13) cable and it is OK again :-) Before it comes back. (0,13) disappeared in run 25805, (0,14) in 25806.
25594 January-24-2002 Physics-60 ! 1249 Physics_emcwf with EMT map 133: Swapped Crate 9, UPC3 with UPC10. (0,13) is still hot, but nothing from UPC3/10. Disconnected and reconnected cables on UPC2. No change. Then much later - in run 25619 - (0,13) also stopped being warm. Looks like we picked up Physics-60.tcf.
25564 January-23-2002 Physics-61 1245 Unmasked the two towers in (theta,phi)=(0,13) and (0.14) as WF run 25544 showed that they were no longer hot. However, they are :-(
25524 January-21-2002 Physics-61 1244 A new EMT tower has just been masked: Theta 0 Phi 14. This tower is adjacent to the tower which was masked last week (theta 0 phi 13) and goes to the same TPB (TPB 7?). Appeared in runs 25521 (possibly), 25522 and 25523.
25400 January-15-2002 Physics-61 1235 Masked off EMT tower in (theta,phi)=(0,13) - EMT map 134. Warm tower, no big effect on the L1 rate, serial number corrupted. Appeared in run 25381.
25280 January-10-2002 Physics-61 1232 Unmasked EMT tower in (theta,phi)=(6,3) - EMT map 133. EMC fiber OK again. Runs 25280, 25281 and 25282 have one bad 'PTD' board.




22047 Aug-20-2001 Physics-57 1132 Unmasked previously hot EMT tower in (theta,phi)=(0,3). Using PHYSICS_EMTTEST for the moment.
21829 Aug-10-2001 Physics-57 1112 New TSF lookup table for the DCT.
21794 Aug-08-2001 Physics-57 1102 Masked off hot EMT tower in (theta,phi)=(0,3).
21050 Jun-27-2001 Physics-57 1052 New EMT configuration.
20315 May-20-2001 Physics-57 1023 Masked off hot EMT tower.
19967 May-03-2001 Physics-57 1017 Unmasked EMT tower (0,14).
19472 Apr-09-2001 Physics-56 998 DCT map that uses rationalized names.
19429 Apr-07-2001 Physics-56 994 Masked off hot EMT tower.
19414 Apr-06-2001 Physics-56 982 New EMT timeAlignment (-3) and a new EMT CyclicTC "object name" without space.
19053 Mar-22-2001 Physics-53 963 New GLT global configuration.
18627 Mar-02-2001 Physics-53 945 Unmasked EMT towers (theta,phi)=(0,13) and (0,14).
18597 Feb-28-2001 Physics-53 942 New TrgConfig with relaxed GLT prescales, random triggers removed from the L3 physics input, and Bhabha prescales for 1-1.5 10^33.
18597 Feb-28-2001  Physics-51 941/942 Unmasked EMT towers (theta,phi)=(1,37) and (2,37).
18522 Feb-24-2001  Physics-52 936 New TrgConfig that gives higher priorities to DCH based lines to aleviate the current EMC timing problems.
18522 Feb-24-2001  Physics-51 935/936 EMT changes.
18398 Feb-19-2001  Physics-51 932 EMT changes.
18397 Feb-19-2001  Physics-51 931 EMT changes.
18328  Feb-16-2001 Physics-51 930 EMT changes.
18245  Feb-13-2001 Physics-51 917 Replacement of UPC firmware PROMs. Only 7 towers masked out.
18190  Feb-12-2001 Physics-51 910 Masking off three more hot EMT towers.
18187  Feb-12-2001 Physics-51 908 Masking off three more hot EMT towers.
18177  Feb-11-2001 Physics-51 905 New TrgConfig for the energy scan.
18149  Feb-11-2001 Physics-44 901 DCT map that includes the new TSF lookup table object.
18145  Feb-11-2001 Physics-44 900 Masking off more hot EMT towers.
18123  Feb-10-2001 Physics-44 899 Masking off more hot towers.
18114  Feb-10-2001 Physics-44 895 To solve the EMT hot tower problem .....
18074  Feb-07-2001 Physics-44 893 Masked off hot EMT tower in (theta,phi)=(5,2).
18053  Feb-06-2001 Physics-44 887 To pick up the fully persistent EMT objects so offline jobs can read the EMT thresholds and parameters.




 17105  Oct-28-2000 Physics-48 858 EMC->New EMT time alignment (-3 to -1)
 17030  Oct-26-2000 Physics-48 854 New GLT config: new 2BM line, 2M&A+, and activating prescaled lines.
 16859  Oct-20-2000 Physics-47 ? Removed DCT FEX truncation for TSF data. Run 16858--16991, between 2 DOCAD tests had inconsistent TSF LUT between TSF and ROM.
 16497  Oct-06-2000 Physics-37 ? EMC->EMT crystal threshold lowered 20->10MeV. Also fixed Hi-E out of time digis since run 16830.
 11465  Feb-07-2000 Physics-37 685 EMT M threshold raised 102->120 MeV.
 11332  Feb-04-2000 Physics-37 684 Further prescaled D2,2M /2; Other prescaled lines by /100 due to high background.




 10046  Oct-26-1999 Physics-all-27 540 Replaced M*&1U with M*&5U (new IFT def) to reduce rate.
The EMC energy scale changed during shutdown, EMT thresholds effctively lower by 20%.
  7789  Aug-14-1999 Physics-all-18 ? Redefined L1 with reduced rate 
a) GLT back-back * object angle cut~120 degrees; 
b) Increased prescales for D2,2M,M*; 3M,M*+D2* prescaled; 
c) Reordered all trigger lines with high mult DCT at top; 
d) Cyclic-1 (1Hz) trigger installed. 
Aug/26: Line 15/16 confusion corrected; GLT match phispan=3 bins.
Also added light pulser trigger=> -all-19 to 21.
Oct/23: Key 534, Run 9948, GLT output line delay realigned.
Oct/24: Run 9992, BLT channel mask on for DCH SL5,6 HV problem.
  5837  Jun-13-1999 Physics-all-13 ? Physics triggers for U(4s) scan. 
a) Remove A' on 3-track trigger (left from 1/3 B field ); 
b) Add 1B,1M for background monitoring; 
c) Replace U3,M*&1B with M*&D2,M*&A+; 
d) Doubled prescale rates for loose open triggers. 
  2660 May-23-1999 Physics-all-0 ? Physics trigger with all DCT+EMT+IFT & B field on. 
  2610
May-20-1999
Physics-emt-2 ? First colliding beam triggers mainly EMT only 
(prelim version started on May/19; May/20 added U=3; May/22 DCT triggers at lines18-24 for testing B=1000)
  2513
May-18-1999
Physics-test ? Test version of beam physics triggers (all DCT+EMT) 
(also did some tests in some runs in last few days)

Level-3 Physics Configuration


Start
Run
Start
Time
Configuration
Name
Key
Comments
17105  Oct-28-2000 Physics-50 858 Activated phi+gamma filter.
17091  Oct-28-2000 Physics-49 857 L3 EMC time window cut installed. Crystal threshold 30->20MeV.
17066  Oct-27-2000 Physics-49 857 Activated improved Bhabha veto.
16859  Oct-20-2000 Physics-47 852 Removed DCT FEX truncation for TSF data. Run 16858--16991, between 2 DOCAD tests had inconsistent TSF LUT between TSF and ROM.
15605  Aug-31-2000 Physics-47 818 2nd phase L3 Dch bias fix & ghost track removal. Also fixed cosmic filter momentum window cut bug.
13446  May-23-2000 Physics-44 767 1st phase L3 Dch bias fix. Bhabha Opr prescale increased. FCGM line 26,27 named as pulser lines and included in L1Open.
12556  Mar-29-2000 Physics-42 742 Removed 1Y1B, 1Y from physics input. Many Opr line prescales increased.
11332  Feb-04-2000 Physics-37 684 Cyclic+bunch OPR output prescaled  further by/10 -> ~0.1Hz each.
11018  Jan-24-2000 Physics-36 638 Took out obsolete mu-pair script.
10971  Jan-22-2000 Physics-35 634  Standard L3 configuration with Bhabha veto after E-scan. 
10865  Jan-16-2000 Physics-33 615 L3 with new EMC clustering &  new EMC filters.
Added Hadron/Bhabha monitoring.
Added output lines with appropriate prescale for OPR.
*special config for E-scan: DCH+EMC bypass Bhabha veto.
10601  Nov-26-1999 Physics-all-27 578 Activate new Bhabha veto/accept-prescale, GamGam, RadBhabha. Also prescale OnlLumi and remove 2E&1Y pass throughs. Release 2.18.5 / 8.3.1-online-2. 
9948  Oct-23-1999 Physics-all-26 534 New Bhabha accept filter in flagging mode & remove old filters.
9600  Sep-26-1999 Physics-all-24 491 New Bhabha veto in flagging mode. Online Lumi monitor added.
8846  Sep-02-1999 Physics-all-22 461 New IPtrack 1+2 filters replaced old 1-track filter.
8444  Aug-25-1999 Physics-all-20 450 GLT global TC in DB & fixed L1 line 15<->16.
8020  Aug-19-1999 Physics-all-19 446 Added light pulser output line (script list changed)
7789  Aug-14-1999 Physics-all-18 425 L1 pass through 1%->0.5%. 
Many L3 configuration script/line changes in July/Aug are not
detailed here: Bhabha scripts, new IPtrk scripts etc.
5837  Jun-13-1999 Physics-all-13 309 The duplicate set of BunchOut lines removed. 
L1 pass through rate 2%->1%. 
1Y pass through replaced by 1Y&1B. 
Cosmic filter d0,z0 tightened . 
2675 May-24-1999 Physics-all-4 206 Pass through 2E & 1Y ; Neutral filter Ecut>2 GeV 
(Physics-all-5 key 207 : L1 pass through 10% used briefly for runs 5256-5257)
2660 May-23-1999 Physics-all-0 199 First colliding beam physics setup

Background Run Configuration


Start
Run
Start
Time
Configuration
Name
Key
Comments
10604  Nov-27-1999 Background-28 583 L1 pass through 3%; Cyclic-1 10Hz; BunchX 8 Hz. Based on Physics-all-27 for L1 & L3. 
8843  Sep-02-1999 Background-21 459 L1 pass through 3%; Cyclic-1 10Hz; BunchX 8 Hz. 
8843  Aug-16-1999 Background-18 426 L1 pass through 1%; Cyclic-1 10Hz; BunchX 8 Hz. 

Cosmic Configuration 

Start 
Run
Start 
Time
Configuration 
name
Comments
52732
Mar-8-2005
0x14ff
Cosmic config with DCZ; Move IFT lines to lower priority (LST/RPC timing split) and IFT lines revised with more options.
51464
Oct-29-2004
0x13fd
Cosmic config with DCZ.
17232
Nov-29-2000
Cosmics-18 Made pure PTD A' lines. (Key 863)
9755
Oct-15-1999
Cosmics-10 New IFR trigger definition. (Key 508=L3; key 509=100% L1 pass; 510=509 & EMC threshold 500KeV). Cosmics-11 is 100% L1 passthrough.
Oct/23: Cosmics-12/13 GLT output line delays=0 to realign.
9080
Sep-07-1999
Cosmics-8 Update Cosmics7 Z cut +-50cm; cyclic out ->0.1Hz. (Key 472)
9069
Sep-06-1999
Cosmics-7 DCH+SVT alignment cosmics with new cosmic L1 list. 
L3 cosmic-2-prong Z cut at +-30cm. (Key 468)
1582
Apr-23-1999
Cosmics-5 BaBar roll on cosmics with all trigger components 
(a slightly different test version used since Apr/21)
1100
Feb-04-1999
Cosmics-3 Feb-Apr/99 cosmics with EMC+IFR only
643
Jan-22-1999  Cosmics-2 Jan/99 cosmic run with all DCT+EMT+IFT triggers